ADNAN MUHAMMED

Grade 12,

Amina thinks of a number and subtracts 5/2 from it. She multiplies the result by 8. The result now obtained is 3 times the same number she thought of. What is the number?

 Amina thinks of a number and subtracts 5/2 from it. She multiplies the result by 8. The result now obtained is 3 times the same number she thought of. What is the number? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the number be x,
According to the question, (x–5/2)×8=3x
⇒8x–40/2=3x
⇒8x–3x=40/2
⇒5x=20
⇒x=4
Thus, the number is 4.

Thanks